Hello
Here is a word problem I am having trouble with.
Bob invests in two different retirement funds. He invests some money in a fund with 3% interest and twice as much money in another fund at 4.5% interest. Find the amount deposited at each rate if the total annual interest income is $3,000.
Here is the equation I have written but not sure if it is correct.
Thanks

Answer by mananth(12270)

Part I 3.00% per annum ------------- x
Part II 4.50% per annum ------------ y
y=2x
Interest----- 3000
3.00% x + 4.50% * 2 x = 3000
0.03 x + 0.045 * 2 x = 3000
0.03 x + 0.09 x = 3000
0.12 x = 3000
/ 0.12
x= 25000 --- Fund A
2 x = 50000 Fund B

CHECK
25000 @ 3.00% = 750
50000 @ 4.50% = 2250
Interest= 3000